OpenStudy (anonymous):

what is v if √(v+3)-1=7?

hero (hero):

√(v+3)-1=7 √(v+3)=8 v+3 = 64 v = 61

hero (hero):

yw :)

OpenStudy (anonymous):

what is v if 90=9√25v?

OpenStudy (anonymous):

v=2

hero (hero):

90=9√25v 10 = √25v 100 = 25v 100/25 = v 4 = v

OpenStudy (anonymous):

when u square u should square v also isn't t?

hero (hero):

if the v is under the square root, you don't square it.

OpenStudy (anonymous):

if whole 25v is under sqrt then u right. i thought only 25 is under sqrt

hero (hero):

janella, is the v inside or outside the root?

hero (hero):

That's what I thought. v = 4

OpenStudy (anonymous):

what is n if √3n=√(4n-1)?

hero (hero):

3n = 4n - 1 n = 1

OpenStudy (anonymous):

what is n if √(2n-88)=√(n/6)?

OpenStudy (anonymous):

2n-88=n/6 12n-528=n 11n=528 n=48

OpenStudy (anonymous):

what is n and x if √(x/10)=√(2n-6)?

OpenStudy (anonymous):

since there are 2 unknowns we need 2 eqns to solve them.

OpenStudy (anonymous):

can u help me to solve that equation?

hero (hero):

√(x/10)=√(2n-6) x/10 = 2n-6 x = (2n-6)/10 n = ((x/10) - 6)/2